the length of a rectangle is 6 inches longer than the width. the perimeter of the rectangle is 40 inches how many in inches long is each side
Answer by checkley79(3341)

L=W+6
P=2L+2W
40=2(W+6)+2W
40=2W+12+2W
40=4W+12
4W=40-12
4W=28
W=28/4
W=7 ANS. FOR THE WIDTH.
L=7+6=13 ANS. FOR THE LENGTH.
PROOF:
40=2*13+2*7
40=26+14
40=40